Betancourt International has operations in Arrakis. The balance sheet for this division in Arrakeen solaris shows assets of 22,000 solaris, debt in the amount of 7,000 solaris, and equity of 15,000 solaris.

a. If the current exchange ratio is 1.32 solaris per dollar, what does the balance sheet look like in dollars? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Balance Sheet
Assets Liabilities
Debt $
Equity
Assets $ Total liabilities & equity $



b. Assume that one year from now the balance sheet in solaris is exactly the same as at the beginning of the year. If the exchange rate is 1.47 solaris per dollar, what does the balance sheet look like in dollars now? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Balance Sheet
Assets Liabilities
Debt $
Equity
Assets $ Total liabilities & equity $



c. Assume that one year from now the balance sheet in solaris is exactly the same as at the beginning of the year. If the exchange rate is 1.17 solaris per dollar, what does the balance sheet look like in dollars now? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Balance Sheet
Assets Liabilities
Debt $
Equity
Assets $ Total liabilities & equity $
Answer #1


solution: Part a Balance sheet Assets=22000/1.32 16,666.67 Debt=7000/1.32 5,303.03 equity=15000/1.32 11,363.64 Total assets 1